OpenStudy (anonymous):

PLEASE HELP ME !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!A cubical box is 729 cubic centimeters. What is the length of one side of the box?the answers are A. 27 cm B. 18 cm C. 9 cm D. 3 cm

OpenStudy (anonymous):

please help me

OpenStudy (jdoe0001):

well, we know that \(\bf 729 cm^3 = Length \times Width \times Height\) so 3 units multiplied will yield 729, let's see what \(\bf \large \sqrt[3]{729}\) would give us?

OpenStudy (anonymous):

243

OpenStudy (jdoe0001):

243? let's see 243 * 243 * 243 \(\bf \ne\) 729

OpenStudy (anonymous):

no it dont

OpenStudy (jdoe0001):

right, so what would be \(\huge \bf \sqrt[3]{729} \ \ ?\)

OpenStudy (jdoe0001):

so, whatever happened to \(\huge \bf \sqrt[3]{729} \ \ ?\)